# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v2

@@ 4 Log Depth Enables Recognizing Regular Languages
Constant-depth transformers cannot recognize regular languages, a natural task closely related to state tracking (Liu et al., 2023; Merrill et al., 2024). Liu et al. (2023, Theorem 1) show that a non-uniform variant of log-depth transformer can recognize regular languages using an associative prefix-sum construction (cf. Hillis and Steele Jr, 1986; Blelloch, 1990). However, it is unclear whether standard transformers can implement their construction (Liu et al., 2023, Page 44) for two key reasons:
@@
    Their construction is non-uniform, in the sense that the parameters depend on the input length and depth of the network. This means that it is unclear whether a single transformer could implement it in a way that generalizes to inputs of any length.
@@
    Their result requires two strong simplifications to the transformer architecture: removing residual connections and assuming specific, nonstandard positional encodings. As discussed in Section 2.1, dealing with residual connections is particularly tricky in universal transformers, requiring proper memory management of cells in the residual stream so that outputs from the previous iteration of a layer do not interfere with a later iteration.
Our result, leveraging Lemma 1 (full proof in Appendix C), addressed both of these weaknesses. A single transformer with fixed parameters can recognize strings of any length; moreover this transformer can have residual connections and does not rely on specific positional encodings:
@@
Let $L$ be a regular language over $\Sigma$ and \$$\$\not\in\Sigma$. Then there exists a $(0,8,9)$-universal transformer with causal masking that, on any string $w\$$, recognizes whether $w\in L$ when unrolled to $\lceil\log_{2}\lvert w\rvert\rceil$ depth.
Theorem 1 thus reveals that running a transformer to $\Theta(\log n)$ depth on inputs of length $n$ unlocks new power compared to a fixed-depth transformer. If we do not care that the construction is uniform across layers, we can simplify the block of 8 layers that determines activeness to 1 layer: we simply hardcode the layer index $\ell$ and use a single transformer layer to compute $i\mod\ell$. Thus, the non-uniform construction results in a family of shallower transformers:
@@
Let $L$ be a regular language over $\Sigma$ and \$$\$\not\in\Sigma$. There exists a family of causally masked transformers $\{T_{n}\}_{n=1}^{\infty}$ where $T_{n}$ has $54\lceil\log_{2}n\rceil+$ 5 layers such that, on any string $w\$$ of length $n$, $T_{n}$ recognizes whether $w\in L$.
These results can be extended beyond regular languages: if a $b$-layer transformer can perform some binary associative operation $\oplus:X\times X\to X$, then one can construct an $\Theta(b\log n)$ layer transformer that computes the iterated version on $n$ values, $x_{1}\oplus x_{2}\oplus\ldots\oplus x_{n}\in X$. One example is iterated matrix multiplication. For matrices from a fixed set (e.g., $k\times k$ boolean matrices), Theorem 1 already shows that this task can be performed. However, if the matrices are not from a fixed set (e.g., matrices over $\mathbb{Z}$ or $\mathbb{Q}$ or whose shape depends on $n$), then it is unclear whether log-depth transformers can solve the binary multiplication problem, and thus whether they can solve the iterated version.
@@
Interestingly, while Theorems 1 and 1.1 are about log-depth transformers, they can be turned around to infer bounds on the input length up to which fixed depth transformers (i.e., depth fixed w.r.t. input length) can recognize regular languages. Specifically, given any regular language $L$ and a fixed $d$, Corollary 1.1 implies that there exists a depth $d$ transformer that can recognize strings $w\in L$ as long as $4\lceil\log_{2}\lvert w\rvert\rceil+5\leq d$, [^5] The inequality holds since Corollary 1.1 generalizes to show $T_{n}$ recognizes all strings of length $m\leq n$. which is satisfied if $4\;(1+\log_{2}\lvert w\rvert)+5\leq d$, i.e., $\lvert w\rvert\leq 2^{(d-9)/4}$:
@@
Let $L$ be a regular language over $\Sigma$ and \$$\$\not\in\Sigma$. For any $d\in\mathbb{N}$, there exists a causally masked $d$-layer transformer that, on any string $w\$$ of length at most $12^{(d-9)/4}+$ 1, recognizes whether $w\in L$.
An analogous result holds for universal (i.e., shared parameter) transformers from Theorem 1.
@@ 5 Log Depth Enables Graph Connectivity
In the graph connectivity problem (also referred to as STCON or the reachability problem), the input is a graph $G$, along with a source vertex $s$ and a target vertex $t$. The task is to determine if $G$ has a path from $s$ to $t$. This is a core problem at the heart of many computational questions in areas as diverse as network security, routing and navigation, chip design, and—perhaps most commonly for language models—multi-step reasoning. This problem is known to be complete for the class of logspace Turing machines (Reingold, 2008; Immerman, 1998), which means that, under common complexity conjectures, it cannot be solved accurately by fixed-depth transformers, which can only solve problems in the smaller class $\mathsf{TC}^{0}$. However, graph connectivity can be expressed by log-depth threshold circuits ($\mathsf{TC}^{1}$, Barrington and Maciel, 2000), which opens up a natural question: Can log-depth transformers, which are in $\mathsf{TC}^{1}$, solve graph connectivity? We prove the following, which shows that the answer is yes (see proof sketch below, full proof in Appendix D):
@@
There exists a $(17,2,1)$-universal transformer $T$ with both causal and unmasked heads that, when unrolled $\lceil\log_{2}n\rceil$ times, solves connectivity on (directed or undirected) graphs over $n$ vertices: given the $n\times n$ adjacency matrix of a graph $G$, $n^{3}$ padding tokens, and $s,t\in\{1,\ldots n\}$ in unary, $T$ checks whether $G$ has a path from vertex $s$ to vertex $t$.
@@
Repeated Layers. The repeated layers alternates between computing the $C_{\ell}$ and the $B_{\ell}$ predicates for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$. The idea is to compute $C_{\ell}(i,k,j)$ in the $n^{3}$ padding tokens by attending to positions $(i,k)$ and $(k,j)$ and retrieving $B_{\ell-1}(i,k)$ and $B_{\ell-1}(k,j)$. Similarly, $B_{\ell}(i,j)$ is computed in the $n^{2}$ input positions via uniform attention over padding positions $(i,k^{\prime},j)$ that store $C_{\ell}(i,k^{\prime},j)$.
Final Layers. Finally, in layer $182\lceil\log n\rceil+$ 18, the final token uses a head that attends with query $\langle\phi(s),\phi(t)\rangle$ corresponding to the source and target nodes $s$ and $t$ mentioned in the input, attending solely to the position with coordinates $(s,t)$, and retrieving the final value $B_{\lceil\log n\rceil}(s,t)$. ∎
@@ 6 Comparing Scaling Depth to Scaling Width or Chain of Thought
Our results show that looping layers enables transformers to solve problems likely outside $\mathsf{TC}^{0}$. We now consider how looping compares in expressive power to other ways to add computation to transformers. Rather than increasing depth by repeating layers, one can increase a transformer’s width via a larger model dimension (Definition 1) or padding tokens (Pfau et al., 2024). Whereas slightly increasing depth likely expands expressive power beyond $\mathsf{TC}^{0}$, we show that achieving expressivity beyond $\mathsf{TC}^{0}$ via depth likely requires superpolynomial in width, which is intractable. In contrast to repeating layers, another way to extend inference-time computation is using chain-of-thought (CoT) steps. We thus compare the expressive power achieved repeated layers with CoT steps.
@@
Our Corollaries 1.2 and 2.1 show that minimally growing a transformer’s depth allows it to express key problems that are likely outside $\mathsf{TC}^{0}$. In contrast, Theorem 3 (which extends Merrill and Sabharwal (2023a); for completeness, Appendix E gives a sketch) shows that, if depth remains fixed, width must increase drastically with sequence length to enable expressive power outside $\mathsf{TC}^{0}$.
@@
Let $T$ be a fixed-depth transformer whose width (model dimension or padding tokens; Pfau et al., 2024) grows as a polynomial in $n$ and whose weights on input length $n$ (to accommodate growing width) are computable in $\mathsf{L}$. Then $T$ can be simulated in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Thus, to solve reasoning problems outside $\mathsf{TC}^{0}$ over a context length $n$, growing depth is much more efficient than growing width. Of course, there may be other types of problems (e.g., those that are knowledge intensive or very parallelizable) where growing width might be more important than growing depth. Petty et al. (2024) provide an interesting empirical investigation of this choice on language modeling, semantic parsing, and other tasks.
@@
Merrill and Sabharwal (2024, Theorem 4) analyze the power of transformers with $\mathrm{O}(\log n)$ CoT steps, showing it is at most $\mathsf{L}$. However, we have shown that transformers with $\Theta(\log n)$ depth can solve directed graph connectivity, which is $\mathsf{NL}$-complete: this suggests growing depth has some power beyond growing CoT unless $\mathsf{L}=\mathsf{NL}$. In fact, the $\mathrm{O}(\log n)$ CoT steps result can be strengthened (Li et al., 2024, Figure 10; for completeness, Appendix E gives a sketch) to an upper bound of $\mathsf{TC}^{0}$:
@@
Thus, while giving a model $\mathrm{O}(\log n)$ CoT steps does not increase its expressive power beyond $\mathsf{TC}^{0}$, our Theorems 1 and 2 allow $\Theta(\log n)$ to solve key problems that are (likely) outside $\mathsf{TC}^{0}$. This demonstrates an advantage of dynamic depth over CoT as a form of inference-time compute for reasoning problems including regular language recognition and graph connectivity. It would be interesting to explore this comparison more generally for other problems.
@@ 7 Experiments: Learning to Recognize Regular Languages
Our theory characterizes the depth and width required to express regular language recognition and graph connectivity. Specifically, Theorem 1 predicts that recognizing regular languages over strings of length $n$ is empirically possible with depth proportional to $\log n$. On the other hand, Theorem 3 predicts that the width would need to scale superpolynomially. Here, we aim to empirically measure how much depth and width transformers require in practice when trained to recognize regular languages. We will find that expressibility and learnability are highly aligned here: transformers with log depth can learn to recognize regular languages, whereas depth must increase exponentially with $n$. Moreover, we can empirically quantify the constant factors in these relationships.
@@
Figure 2: Strong linear fits imply theory/experiment match for modeling the impact of depth (left, $d=4.8\log_{2}n-15.8$ with $r^{2}=0.93$) and width (right, $n=7.2\log_{2}w-41.7$ with $r^{2}=0.98$) on effective context length for the $A_{5}$ state tracking task, a canonical hard regular language recognition problem. As predicted by Theorems 1 and 3, to recognize strings of length $n$, depth only needs to increase minimally $\propto\log n$ while width must increase drastically as $\exp(\Theta(n))$.
@@
We report on an extensive set of experiments to address these questions, training models of different depths and widths on the $A_{5}$ state tracking task (Merrill et al., 2024), which is a canonical testbed for hard regular language recognition (Theorem 1). The input to the task is a sequence of elements in $A_{5}$ (the group of even permutations over 5 elements), and the label at each token is the cumulative product of previous permutations up to and including that token (which is itself an element of $A_{5}$).
@@
We train several (non-universal) transformers with the same architecture used by Merrill et al. (2024) on  100 million $A_{5}$ sequences of varying lengths up to 1024. To understand the impact of depth and width in a controlled way, we train two series of transformers: the first with width fixed to 512 and depth varying in $\{6,9,12,15,18,21,24\}$, and the second with depth fixed to 6 and width varying in $\{128,256,512,1024\}$. See Appendix F for further details about our training procedure. After each model is trained, we measure accuracy at each token index from 1 to 1024 and define $n^{*}$ as the maximum token index at which the model achieved at least 95% validation accuracy. As we trained several seeds with the same depth and width, we aggregate these results across all models with the same depth and width by taking the best-performing (max $n^{*}$) model. We then plot $n^{*}$, which represents the effective context length up to which a model can solve the $A_{5}$ problem, as a function of either depth or width, holding the other variable fixed. We then evaluate if the predicted theoretical relationships between depth, width, and context length hold via an $r^{2}$ statistic.
@@
The results are shown in Figure 2. When varying depth (left plot), there is a very strong positive correlation ($r^{2}=0.93$) between model depth (x-axis) and $\log n^{*}$ (y-axis, log scale), the effective (log) context length till which it can solve problems with high accuracy. When varying width (right plot) there is an even stronger positive correlation ($r^{2}=0.98$) between log width (x-axis, log scale) and $n^{*}$ (y-axis). These results provide strong empirical support for our theoretical predictions that, to recognize regular languages over strings of length $n$, increasing depth logarithmically in $n$ will suffice (Theorem 1), but width must increase exponentially in $n$ (Theorem 3). Figure 2 also gives us a strongly predictive functional form to quantify the impact of scaling depth or width on the effective context length for regular language recognition. The empirical slope for the depth relationship is is 4.8 layers per log tokens. This is less than the slope of 8 derived for universal transformers in Theorem 1, but slightly greater than the theoretical coefficient of 4 for transformers whose depth grows non-uniformly with context length. Thus, our transformers have learned a construction whose depth coefficient is comparable to what we showed was possible in theory, though perhaps slightly more wasteful than it needs to be. Overall, these empirical results show that, in practice, the impact of depth and width on effective context length for regular language recognition aligns with our theoretical predictions, and we are able to empirically fit the quantitative coefficients in the relationships.

### v3

@@ 4 Log Depth Enables Recognizing Regular Languages
Constant-depth transformers cannot recognize regular languages, a natural task closely related to state tracking (liu2023transformers; merrill2024illusion). liu2023transformers555saunshi2025reasoning also give a log-depth transformer construction for the regular language recognition problem. It is, however, not fully uniform as positions are encoded with vectors of length $\mathrm{O}(\log n)$. show that a variant of log-depth transformers can recognize regular languages using an associative prefix-sum construction (cf. hillis1986data; BlellochTR90). However, it is, prima facie, unclear whether the fully uniform model of transformers we study—where the parameters cannot change at all with nn—can implement their construction for two key reasons:
@@
    To handle information routing, liu2023transformers require parameters that are not fully uniform, meaning they can depend on the input length and depth. This leaves it unclear whether a single transformer (with a fixed set of parameters) could solve the task across all input lengths. In other words, their work leaves it unclear whether a single transformer could implement the approach in a way that generalizes to inputs of any length.
@@
    liu2023transformers also make several simplifications to the transformer architecture: they add non-standard positional embeddings and remove residual connections and layer-norm. While one could adapt their construction to handle residual connections, it is not clear how to do this while also making their construction uniform, which requires proper memory management of cells in the residual stream (Section˜2.1).
Our result, using Lemma˜1, addresses both of these weaknesses. It shows, for the first time, that a single transformer with fixed parameters (w.r.t. input length $n$) can recognize strings of any length; moreover this transformer does not require specific positional encodings, allows for layer-norm (in fact leverages it), and allows for residual connections while remaining fully uniform.
@@
Let $L$ be a regular language over $\Sigma$ recognized by a (non-)deterministic finite automaton with states $Q$. Let $\mathdollar\not\in\Sigma$. Then there exists a causally masked $(0,8,9)$-universal transformer with
[…]
that, on any string $w\mathdollar$, recognizes whether $w\in L$ when unrolled to $\lceil\log_{2}\lvert w\rvert\rceil$ depth.
Proof in Appendix˜C. Theorem˜1 reveals that running a transformer to $\Theta(\log n)$ depth on inputs of length $n$ unlocks new power compared to a fixed-depth transformer, assuming $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$. If we do not care that the construction is uniform across layers, we can simplify 8-layer block that determines activeness to 1 layer: we simply hardcode the layer index $\ell$ and use a single transformer layer to compute $i\mod\ell$. Thus, the non-uniform construction results in a shallower transformer family:
@@
Let $L$ be a regular language over $\Sigma$ and $\mathdollar\not\in\Sigma$. There exists a family of causally masked transformers $\{T_{n}\}_{n=1}^{\infty}$ where $T_{n}$ has $4\lceil\log_{2}n\rceil+5$ layers such that, on any string $w\mathdollar$ of length $n$, $T_{n}$ recognizes whether $w\in L$.
These results can be extended beyond regular languages: if a $b$-layer transformer can perform some binary associative operation $\oplus:X\times X\to X$, then one can construct an $\Theta(b\log n)$ layer transformer that computes the iterated version on $n$ values, $x_{1}\oplus x_{2}\oplus\ldots\oplus x_{n}\in X$. One example is iterated matrix multiplication. For matrices from a fixed set (e.g., $k\times k$ boolean matrices), Theorem˜1 already shows that this task can be performed. However, if the matrices are not from a fixed set (e.g., matrices over $\mathbb{Z}$ or $\mathbb{Q}$ or whose shape depends on $n$), then it is unclear whether log-depth transformers can solve the binary multiplication problem, and thus whether they can solve the iterated version.
@@
Interestingly, while Theorems˜1 and 1.1 are about log-depth transformers, they can be turned around to infer bounds on the input length up to which fixed depth transformers (i.e., depth fixed w.r.t. input length) can recognize regular languages. Specifically, given any regular language $L$ and a fixed $d$, Corollary˜1.1 implies that there exists a depth $d$ transformer that can recognize strings $w\in L$ as long as $4\lceil\log_{2}\lvert w\rvert\rceil+5\leq d$, [^6] The inequality holds since Corollary 1.1 generalizes to show $T_{n}$ recognizes all strings of length $m\leq n$. which is satisfied if $4\;(1+\log_{2}\lvert w\rvert)+5\leq d$, i.e., $\lvert w\rvert\leq 2^{(d-9)/4}$:
@@
Let $L$ be a regular language over $\Sigma$ and $\mathdollar\not\in\Sigma$. For any $d\in\mathbb{N}$, there exists a causally masked $d$-layer transformer that, on any string $w\mathdollar$ of length at most $2^{(d-9)/4}+1$, recognizes whether $w\in L$.
An analogous result holds for universal (i.e., shared parameter) transformers from Theorem˜1.
@@ 5 Log Depth Enables Graph Connectivity
In the graph connectivity problem (also referred to as STCON or the reachability problem), the input is a graph $G$, along with a source vertex $s$ and a target vertex $t$. The task is to determine if $G$ has a path from $s$ to $t$. This is a core problem at the heart of many computational questions in areas as diverse as network security, routing and navigation, chip design, and—perhaps most commonly for language models—multi-step reasoning. This problem is known to be complete for the class of logspace Turing machines (rheingold2008undirected; immerman1998descriptive), which means that, under common complexity conjectures, it cannot be solved accurately by fixed-depth transformers, which can only solve problems in the smaller class $\mathsf{TC}^{0}$. However, graph connectivity can be expressed by log-depth threshold circuits ($\mathsf{TC}^{1}$, barrington2000lecture), which opens up a natural question: Can log-depth transformers, which are in $\mathsf{TC}^{1}$, solve graph connectivity?
sanford2024understanding provide results showing that a non-uniform log-depth transformer with arbitrarily powerful feedforward nets can solve variants of graph connectivity. We prove that even a fully uniform log-depth transformer can solve graph connectivity (proof sketch below, full proof in Appendix˜D):
@@
There exists a $(17,2,1)$-universal transformer $T$ with both causal and unmasked heads, fixed model dimension $m$, and fixed feedforward width $w$ that, when unrolled $\lceil\log_{2}n\rceil$ times, solves connectivity on (directed or undirected) graphs over $n$ vertices: given the $n\times n$ adjacency matrix of a graph $G$, $n^{3}$ padding tokens, and $s,t\in\{1,\ldots n\}$ in unary, $T$ checks whether $G$ has a path from vertex $s$ to vertex $t$.
@@
Repeated Layers. The repeated layers alternate between computing the $C_{\ell}$ and the $B_{\ell}$ predicates for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$. The idea is to compute $C_{\ell}(i,k,j)$ in the $n^{3}$ padding tokens by attending to positions $(i,k)$ and $(k,j)$ and retrieving $B_{\ell-1}(i,k)$ and $B_{\ell-1}(k,j)$. Similarly, $B_{\ell}(i,j)$ is computed in the $n^{2}$ input positions via uniform attention over padding positions $(i,k^{\prime},j)$ that store $C_{\ell}(i,k^{\prime},j)$.
Final Layers. Finally, in layer $2\lceil\log n\rceil+18$, the final token uses a head that attends with query $\langle\phi(s),\phi(t)\rangle$ corresponding to the source and target nodes $s$ and $t$ mentioned in the input, attending solely to the position with coordinates $(s,t)$, and retrieving the final value $B_{\lceil\log n\rceil}(s,t)$. ∎
@@ 6 Comparing Scaling Depth to Scaling Width or Chain of Thought
Our results show that looping layers enables transformers to solve problems likely (conditionally under the conjecture that $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$) outside $\mathsf{TC}^{0}$. We now consider how looping compares in expressive power to other ways to add computation to transformers. Rather than increasing depth by repeating layers, one can increase a transformer’s width via a larger model dimension (Definition˜1) or padding tokens (pfau2024lets). Whereas slightly increasing depth likely expands expressive power beyond $\mathsf{TC}^{0}$, we show that achieving expressivity beyond $\mathsf{TC}^{0}$ via width likely requires superpolynomial width, which is intractable. In contrast to repeating layers, another way to extend inference-time computation is using chain-of-thought (CoT) steps. We thus compare the expressive power achieved repeated layers with CoT steps.
@@
Our Corollaries˜1.2 and 2.1 show that minimally growing a transformer’s depth allows it to express key problems that are likely outside $\mathsf{TC}^{0}$. In contrast, Theorem˜3 (which extends merrill-sabharwal-2023-parallelism; for completeness, Appendix˜E gives a sketch) shows that, if depth remains fixed, width must increase drastically with sequence length to enable expressive power outside $\mathsf{TC}^{0}$.
@@
Let $T$ be a fixed-depth transformer whose width (model dimension or padding tokens; pfau2024lets) grows at most polynomial in $n$ and whose weights on input length $n$ (to accommodate growing width) are computable in $\mathsf{L}$. Then $T$ can be simulated in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Thus, to solve reasoning problems outside $\mathsf{TC}^{0}$ over a context length $n$, growing depth is much more efficient than growing width. Of course, there may be other types of problems (e.g., those that are knowledge intensive or very parallelizable) where growing width might be more important than growing depth. petty-etal-2024-impact provide an interesting empirical investigation of this choice on language modeling, semantic parsing, and other tasks.
@@
merrill-sabharwal-2024-cot analyze the power of transformers with $\mathrm{O}(\log n)$ CoT steps, showing it is at most $\mathsf{L}$. However, we have shown that transformers with $\Theta(\log n)$ depth can solve directed graph connectivity, which is $\mathsf{NL}$-complete: this suggests growing depth has some power beyond growing CoT unless $\mathsf{L}=\mathsf{NL}$. In fact, the $\mathrm{O}(\log n)$ CoT steps result can be strengthened (li2024chain, Figure 10; for completeness, Appendix˜E gives a sketch) to an upper bound of $\mathsf{TC}^{0}$:
@@
Thus, while giving a model $\mathrm{O}(\log n)$ CoT steps does not increase its expressive power beyond $\mathsf{TC}^{0}$, our Theorems˜1 and 2 allow $\Theta(\log n)$ to solve key problems that are (likely) outside $\mathsf{TC}^{0}$. This demonstrates an advantage of dynamic depth over CoT as a form of inference-time compute for reasoning problems including regular language recognition and graph connectivity. It would be interesting to explore this comparison more generally for other problems.
@@ 7 Experiments: Learning to Recognize Regular Languages
Our theory characterizes the depth and width required to express regular language recognition and graph connectivity. Specifically, Theorem˜1 predicts that recognizing regular languages over strings of length $n$ is empirically possible with depth proportional to $\log n$. On the other hand, Theorem˜3 predicts that the width would need to scale superpolynomially. Here, we aim to empirically measure how much depth and width transformers require in practice when trained to recognize regular languages. We will find that expressibility and learnability are highly aligned here: transformers with log depth can learn to recognize regular languages, whereas width must increase superpolynomially with $n$. Moreover, we can empirically quantify the constant factors in these relationships.
@@
Figure 2: Strong linear fits imply theory/experiment match for modeling the impact of depth (left, $d=4.8\log_{2}n-15.8$ with $r^{2}=0.93$) and width (right, $n=7.2\log_{2}w-41.7$ with $r^{2}=0.98$) on effective context length for the $A_{5}$ state tracking task, a canonical hard regular language recognition problem. As predicted by Theorems˜1 and 3, to recognize strings of length $n$, depth only needs to increase minimally $\propto\log n$ while width must increase drastically as $\exp(\Theta(n))$.
@@
We report on an extensive set of experiments to address these questions, training models of different depths and widths on the $A_{5}$ state tracking task (merrill2024illusion), which is a canonical testbed for hard regular language recognition (Theorem˜1). The input to the task is a sequence of elements in $A_{5}$ (the group of even permutations over 5 elements), and the label at each token is the cumulative product of previous permutations up to and including that token (which is itself an element of $A_{5}$).
@@
We train several (non-universal) transformers with the same architecture used by merrill2024illusion on  100 million $A_{5}$ sequences of varying lengths up to 1024. To understand the impact of depth and width in a controlled way, we train two series of transformers: the first with width fixed to 512 and depth varying in $\{6,9,12,15,18,21,24\}$, and the second with depth fixed to 6 and width varying in $\{128,256,512,1024\}$. See Appendix˜F for further details about our training procedure. After each model is trained, we measure accuracy at each token index from 1 to 1024 and define $n^{*}$ as the maximum token index at which the model achieved at least 95% validation accuracy. As we trained several seeds with the same depth and width, we aggregate these results across all models with the same depth and width by taking the best-performing (max $n^{*}$) model. We then plot $n^{*}$, which represents the effective context length up to which a model can solve the $A_{5}$ problem, as a function of either depth or width, holding the other variable fixed. We then evaluate if the predicted theoretical relationships between depth, width, and context length hold via an $r^{2}$ statistic.
@@
The results are shown in Figure˜2. When varying depth (left plot), there is a very strong positive correlation ($r^{2}=0.93$) between model depth (x-axis) and $\log n^{*}$ (y-axis, log scale), the effective (log) context length till which it can solve problems with high accuracy. When varying width (right plot) there is an even stronger positive correlation ($r^{2}=0.98$) between log width (x-axis, log scale) and $n^{*}$ (y-axis). These results provide strong empirical support for our theoretical predictions that, to recognize regular languages over strings of length $n$, increasing depth logarithmically in $n$ will suffice (Theorem˜1), but width must increase exponentially in $n$ (Theorem˜3). Figure˜2 also gives us a strongly predictive functional form to quantify the impact of scaling depth or width on the effective context length for regular language recognition. The empirical slope for the depth relationship is is 4.8 layers per log tokens. This is less than the slope of 8 derived for universal transformers in Theorem˜1, but slightly greater than the theoretical coefficient of 4 for transformers whose depth grows non-uniformly with context length. Thus, our transformers have learned a construction whose depth coefficient is comparable to what we showed was possible in theory, though perhaps slightly more wasteful than it needs to be. Overall, these empirical results show that, in practice, the impact of depth and width on effective context length for regular language recognition aligns with our theoretical predictions, and we are able to empirically fit the quantitative coefficients in the relationships.